# Large Multimodal Models for Low-Resource Languages: A Survey

### v1

@@ Abstract
In this survey, we systematically analyze techniques used to adapt large multimodal models (LMMs) for low-resource (LR) languages, examining approaches ranging from visual enhancement and data creation to cross-modal transfer and fusion strategies. Through a comprehensive analysis of 106 studies across 75 LR languages, we identify key patterns in how researchers tackle the challenges of limited data and computational resources. We find that visual information often serves as a crucial bridge for improving model performance in LR settings, though significant challenges remain in areas such as hallucination mitigation and computational efficiency. We aim to provide researchers with a clear understanding of current approaches and remaining challenges in making LMMs more accessible to speakers of LR (understudied) languages. We complement our survey with an open-source repository available at: https://github.com/marianlupascu/LMM4LRL-Survey.
@@ 1 Introduction
Recent advancements in large multimodal models (LMMs) showcased remarkable capabilities in processing and understanding diverse types of data, including text, images, audio and video. Models like GPT-4V, KOSMOS-1 Huang et al. (2023) and PaLM-E Driess et al. (2023) achieved impressive performance levels across various multimodal tasks through their ability to simultaneously process and reason about multiple modalities. However, these developments have primarily focused on high-resource languages, particularly English, leaving a significant gap in supporting the world’s many low-resource languages. The distinction between high-resource (HR) and low-resource (LR) languages is primarily determined by the availability of digital resources and training data. High-resource languages, such as English, Mandarin, and Spanish, benefit from extensive digital corpora, parallel texts, and annotated datasets. In contrast, low-resource or understudied languages, which constitute the majority of the world’s languages, lack sufficient digital resources, standardized datasets, and computational tools. This disparity is particularly pronounced in multimodal contexts, where the scarcity of paired data across modalities (e.g. image-text pairs, audio-text alignments) poses additional challenges.
@@
The motivation for developing multimodal capabilities for LR languages is compelling. First, multimodal processing better reflects how humans naturally communicate and understand information through multiple sensory channels. Second, visual and audio cues can provide crucial contextual information that helps to overcome the limitations of scarce textual data. Third, many LR languages are primarily spoken rather than written, making multimodal approaches particularly relevant for their digital preservation and processing. However, developing multimodal systems for LR languages faces several significant challenges, including: (1) the scarcity of high-quality multimodal datasets in these languages, (2) the lack of standardized evaluation benchmarks, (3) the computational cost of training large-scale models with limited resources, and (4) the complexity of handling different writing systems, dialects, and cultural contexts. Moreover, the problem of catastrophic forgetting when adapting pre-trained models to new languages and the challenge of maintaining performance across different modalities pose significant technical hurdles.
@@
Figure 2: Distribution of papers across 75 low-resource languages, representing 106 papers. Hindi leads with 26 studies, followed by Arabic (18), Bengali and Malayalam (16 each), and Tamil (11). The remaining languages have less than 5 papers each. The data highlights the disparity in research focus among LR languages, with a few languages receiving more focus, while many others remain understudied in the context of multimodal learning. Some papers simultaneously address multiple languages, contributing to the individual language counts. HR languages such as English, Chinese, Mandarin and Spanish are excluded from this chart.
@@
In this context, we survey research articles from 2018 to the end of 2024 that specifically study LMMs for LR languages. We focus on works that go beyond simple cross-lingual transfer or translation, examining techniques that leverage multiple modalities to improve model performance. Our analysis reveals several interesting patterns in how researchers approached multimodal learning for LR languages. As shown in Figure 1, text-image combinations dominate the research landscape, appearing in 67 papers (63% of surveyed works), while more complex combinations incorporating audio and video remain less explored. In addition, the distribution of research focus across languages is notably uneven, as illustrated in Figure 2, with Hindi (26 papers), Arabic (18 papers) and Bengali (16 papers) receiving significant focus, whereas 40 other languages are each represented by a single study. This disparity highlights the need for a broader coverage of understudied languages in multimodal research.
@@
Some recent surveys have explored various aspects of multimodal language models. Yin et al. [2024] provide a comprehensive overview of LMM architectures, training strategies, and applications, while Wang et al. [2023a] focus on pre-training techniques and model architectures. However, these works primarily focus on high-resource languages, and do not specifically address the unique challenges and solutions for LR languages in multimodal contexts. To the best of our knowledge, our survey is the first to address multimodal learning for understudied languages.
@@
  • We provide the first comprehensive analysis of LMMs specifically focused on LR languages, examining 106 studies across 75 languages.
  • We develop a novel taxonomy (see Figure 3) that categorizes existing approaches into six main categories: data creation and engineering, fusion techniques, visual enhancement, cross-modal transfer, synthetic data generation, and architectural innovations.
@@ 2 Taxonomy
In Figure 3, we systematically organize LMMs for LR languages into six main categories, reflecting both the current state of the field and the primary research strategies for addressing challenges in the context of under-represented languages. The first two categories focus on constructing high-quality resources. While the first category discusses multimodal data creation either from scratch or via extending existing datasets, the second approach centers on synthetic data generation, which automatically expands available resources via back-translation and image-based generation. Building upon this work, we present several multimodal fusion techniques and provide various strategies for effectively combining this information, ranging from early and late fusion to more complex hybrid approaches. In the fourth category, we illustrate visual enhancement techniques that harness visual information through image-guided translation and visual disambiguation methods, highlighting their importance for improving translation quality and resolving ambiguities. Expanding from the single-modality solutions, the next category focuses on cross-modal transfer learning approaches that can facilitate knowledge sharing based on both modality transfer and language transfer. Finally, our last category comprises architectural innovations specifically tailored for multimodal tasks in the context of LR languages. We structure the remainder of this article according to our novel taxonomy shown in Figure 3.
@@ 3 Multimodal Data Creation
Dataset creation from scratch. Dataset creation from scratch has emerged as a crucial approach for enabling multimodal research in LR languages, particularly for sentiment analysis and specific language tasks. Multiple research teams have focused on creating specialized datasets through direct data collection and annotation, such as collecting Arabic videos with multimodal features for sentiment analysis Najadat and Abushaqra (2018), building comprehensive Tamil and Malayalam video review datasets Chakravarthi et al. (2021), and developing new corpora for languages such as Malay Taylor and Fauzi (2024). A significant trend has been the creation of meme-based datasets, with efforts focused on Bengali, through MemoSen and MUTE Hossain et al. (2022a, b), and Romanian, through RoMemes Păiş et al. (2024), all incorporating multiple levels of annotation.
@@
These dataset creation efforts have expanded beyond sentiment analysis to encompass other crucial applications, such as sign language recognition with ArabSign Luqman (2023) and multi-purpose datasets like BIG-C for Bemba Sikasote et al. (2023) and Arabic multimodal sentiment collections Haouhat et al. (2023). These projects typically involve careful quality control by using multiple annotators, standardized recording environments, and expert validation, demonstrating a shift toward building comprehensive resources specifically designed for LR languages, rather than relying on translation or transfer from high-resource languages.
@@
Dataset extension. In addition to building data from scratch in the context of LR language understanding, there have been several efforts for leveraging existing datasets of rich-resource languages and building upon them. Sen et al. [2022] introduced the Bengali Visual Genome (BVG) dataset, which extends the Visual Genome dataset Krishna et al. (2017) with Bengali translations and annotations, enabling the development and evaluation of multimodal models for Bengali-English machine translation (MT) and image captioning. Similarly, Abdulmumin et al. [2022] created the Hausa Visual Genome (HaVG) dataset by translating a subset of the Visual Genome dataset into Hausa, providing a valuable resource for English-to-Hausa multimodal MT. Building upon prior work and continuing the focus on the Hausa language, Parida et al. [2023] introduced the Hausa Visual Question Answering (HaVQA) dataset, which adapts question-answer pairs from the Visual Genome dataset to the Hausa language through manual translation, creating the first visual question-answering (VQA) dataset for Hausa.
@@
Apart from the focus on African languages, Saichyshyna et al. [2023] extended the Multi30K dataset Elliott et al. (2016) to include Ukrainian translations and captions, facilitating integrated vision and language research in Ukrainian. More recently, Lovenia et al. [2024] presented SEACrowd, a comprehensive multilingual and multimodal data hub and benchmark suite for Southeast Asian languages, which covers 13 tasks across three modalities (text, image, and audio) and 38 Southeast Asian indigenous languages. These tasks include MT, image captioning, VQA, and speech recognition. To build SEACrowd, the authors leveraged already existing datasets, but they also created new data through crowd-sourcing and manual annotation.

### v2

@@ Abstract
In this survey, we systematically analyze techniques used to adapt large multimodal models (LMMs) for low-resource (LR) languages, examining approaches ranging from visual enhancement and data creation to cross-modal transfer and fusion strategies. Through a comprehensive analysis of 117 studies across 96 LR languages, we identify key patterns in how researchers tackle the challenges of limited data and computational resources. We categorize works into resource-oriented and method-oriented contributions, further dividing contributions into relevant sub-categories. We compare method-oriented contributions in terms of performance and efficiency, discussing benefits and limitations of representative studies. We find that visual information often serves as a crucial bridge for improving model performance in LR settings, though significant challenges remain in areas such as hallucination mitigation and computational efficiency. In summary, we provide researchers with a clear understanding of current approaches and remaining challenges in making LMMs more accessible to speakers of LR (understudied) languages. We complement our survey with an open-source repository available at: https://github.com/marianlupascu/LMM4LRL-Survey.
journal: Information Fusion
[…]
[unibuc]organization=Department of Computer Science, University of Bucharest,country=Romania
@@ 1 Introduction
Recent advancements in large multimodal models (LMMs) showcased remarkable capabilities in processing and understanding diverse types of data, including text, images, audio and video. Models like GPT-4V, KOSMOS-1 [NEURIPS2023_e425b75b] and PaLM-E [driess2023palm] achieved impressive performance levels across various multimodal tasks through their ability to simultaneously process and reason about multiple modalities. However, these developments have primarily focused on high-resource languages, particularly English, leaving a significant gap in supporting the world’s many low-resource languages. The distinction between high-resource (HR) and low-resource (LR) languages is primarily determined by the availability of digital resources and training data. High-resource languages, such as English, Mandarin, and Spanish, benefit from extensive digital corpora, parallel texts, and annotated datasets. In contrast, low-resource or understudied languages, which constitute the majority of the world’s languages, lack sufficient digital resources, standardized datasets, and computational tools. This disparity is particularly pronounced in multimodal contexts, where the scarcity of paired data across modalities (e.g. image-text pairs, audio-text alignments) poses additional challenges.
A recent analysis [dotan2024invisible] identified $27\%$ of languages as “Invisible Giants”, i.e. demographically robust yet digitally absent, highlighting that resource scarcity is institutionally constructed rather than inherent. This distinction has practical implications, e.g. LMM development that treats data scarcity merely as technical risks can perpetuate the structural inequalities it ostensibly addresses. We therefore situate our analysis within the UNESCO International Decade of Indigenous Languages (2022-2032) and the CARE principles for Indigenous data governance [carroll2020care], which emphasize community authority over linguistic data. Indeed, the very terminology “low-resource” has been critiqued as colonial and Eurocentric, obscuring the political decisions that produced linguistic marginalization [bird2022decolonising].
The motivation for developing multimodal capabilities for LR languages is compelling. First, multimodal processing better reflects how humans naturally communicate and understand information through multiple sensory channels. Second, visual and audio cues can provide crucial contextual information that helps to overcome the limitations of scarce textual data. Third, many LR languages are primarily spoken rather than written, making multimodal approaches particularly relevant for their digital preservation and processing. However, developing multimodal systems for LR languages faces several significant challenges, including: (1) the scarcity of high-quality multimodal datasets in these languages, (2) the lack of standardized evaluation benchmarks, (3) the computational cost of training large-scale models with limited resources, and (4) the complexity of handling different writing systems, dialects, and cultural contexts. Moreover, the problem of catastrophic forgetting [kirkpatrick2017overcoming] when adapting pre-trained models to new languages and the challenge of maintaining performance across different modalities pose significant technical hurdles.
Literature selection process. We survey research articles from 2018 to 2025 that specifically study LMMs for LR languages. We begin our analysis with 2018 because one of the first large language models (LLMs), BERT [devlin-etal-2019-bert], was introduced that year, marking a significant turning point in the development of modern language modeling techniques. We focus on works that go beyond simple cross-lingual transfer or translation, examining techniques that leverage multiple modalities to improve model performance.
@@
We queried a broad set of digital libraries to ensure representative coverage: ACM Digital Library, IEEE Xplore, ACL Anthology, arXiv, SpringerLink, ScienceDirect, and Google Scholar. During this search, we specifically targeted venues known for frequent LR or multimodal contributions (e.g. ACL, EMNLP, NAACL, COLING, LREC, EACL, WMT, CVPR/ICCV/ECCV workshops, and INTERSPEECH) to ensure that relevant conference and workshop publications are captured. For each of these digital libraries, we formulated several keyword combinations capturing (i) multimodality, (ii) low-resource aspects, and (iii) language or task types. To further improve coverage, we performed a backward and forward search to identify additional relevant work from the reference lists of included papers, and we used citation links to identify more recent follow-up studies.
@@
Figure 2: Distribution of papers across 96 low-resource languages, representing 117 papers. Hindi leads with 31 studies, followed by Arabic (23), Bengali (21), Malayalam (19), Tamil (14), Korean and Yoruba (with 10 papers each). The remaining languages have less than 10 papers each. Languages with only one paper (42 languages) are listed using ISO 639-1 codes. The data highlights the disparity in research focus among LR languages, with a few languages receiving more focus, while many others remain understudied in the context of multimodal learning. Some papers simultaneously address multiple languages, contributing to the individual language counts. HR languages such as English, Chinese, Mandarin and Spanish are excluded from this chart.
Finally, we merged all retrieved records and applied a manual two-stage screening process. We began by reviewing titles and abstracts to remove clearly irrelevant work (e.g. single modality studies or studies exclusively targeting high-resource languages such as English, Mandarin, or Spanish). We then examined the main contributions of the remaining papers to assess whether they were a suitable fit for our survey. Ultimately, a study was included in our survey if it matched all of the following criteria: (a) is multimodal (at least two input modalities), (b) focuses on low-resource languages (at least one of the targeted languages was an underrepresented language), and (c) proposes, adapts or evaluates a multimodal model.
@@
Research focus distribution across LR languages. Our survey reveals several interesting patterns in how researchers approached multimodal learning for LR languages. As shown in Figure 1, text-image combinations dominate the research landscape, appearing in 76 papers (65% of surveyed works), while more complex combinations incorporating audio and video remain less explored. In addition, the distribution of research focus across languages is notably uneven, as illustrated in Figure 2, with Hindi (31 papers), Arabic (23) and Bengali (21 papers) receiving significant focus, whereas 42 other languages are each represented by a single study. On the one hand, this striking disparity highlights the need for a broader coverage of understudied languages in multimodal research. On the other hand, it also warrants critical examination of the factors influencing the distribution of research studies across languages.
Table 1: Factors explaining research disparity across low-resource languages in multimodal NLP. We categorize languages from our survey by paper count and analyze contributing factors.
Factor High Coverage (10+ papers) Medium Coverage (2-9 papers) Low Coverage (1 paper) Institutional capacity Strong local NLP communities (India, Middle East) Emerging research groups Minimal local infrastructure Speaker population Variable (38M–600M) Variable (2M–200M) Often <<1M, but not always Existing resources Multiple benchmarks, pre-trained models Some datasets available Little to no digital presence Script accessibility Shared with HR languages or well-supported Moderate tool support Often unique / unsupported scripts Geographic location Regions with NLP venues (Asia, Middle East) Mixed Often Sub-Saharan Africa, Oceania Geopolitical interest Strategic priority (Arabic post-9/11, Hindi for US-India relations, Korean for East Asia security) Emerging strategic relevance (Ukrainian post-2022, Turkish for NATO relations) No perceived strategic value; excluded from defense / intelligence funding Example languages Hindi, Arabic, Bengali, Malayalam Swahili, Romanian, Turkish, Yoruba Luo, Xhosa, Occitan, Maori
We identify six interacting factors that explain this disparity (see Table 1). First, institutional research capacity plays a dominant role: Rungta et al. [rungta2022geographic] demonstrated that NLP publications are heavily concentrated in North America, Western Europe, and China, with minimal representation from Africa and South America. Languages spoken in regions with established NLP research communities (e.g. Hindi in India, Arabic in the Middle East) benefit from existing infrastructure, funding, and researcher networks. Second, speaker population shows surprising variability: while one might expect larger speaker populations to attract more research, this correlation is weak. For instance, Swahili has approximately 200 million speakers, yet remains underrepresented compared with Malayalam (38 million speakers, 19 papers). Third, digital resource availability creates a self-reinforcing cycle: languages with existing datasets attract more research, which produces more datasets [joshi2020state]. Ranathunga et al. [ranathunga2022languages] showed that even within the same resource class [joshi2020state], languages from higher-GDP regions receive disproportionately more research attention. Fourth, script and typological proximity to high-resource languages facilitates transfer learning research, e.g. Hindi benefits from shared Devanagari script resources, while languages with unique scripts (e.g. Ge’ez for Amharic) face additional barriers. Fifth, geographic location determines access to NLP venues and research networks: languages spoken in regions hosting major conferences (Asia, Middle East, Europe) receive more attention than those in Sub-Saharan Africa or Oceania. Sixth, geopolitical interest drives strategic investment: Arabic NLP surged post-9/11, Ukrainian gained attention after 2022, and US-China AI competition benefits research on Mandarin Chinese, but not on minority languages within China.
These factors have critical implications for researchers working on truly underrepresented languages. The 42 languages with single studies in our survey face a “cold start” problem: without existing benchmarks, baselines, or community momentum, new contributions are harder to contextualize and evaluate [caines2019geographic]. We observe that 88.4% of the world’s languages (Class 0 defined by Joshi et al. [joshi2020state]) have no representation in standard NLP resources whatsoever [joshi2020state]. For researchers targeting these languages, we recommend: (1) prioritizing dataset creation with community involvement over model development, (2) leveraging typologically similar languages for transfer rather than defaulting to English, and (3) publishing in venues with explicit low-resource tracks (e.g. AfricaNLP, AmericasNLP, etc.) to build critical mass within language-specific research communities.
Research investment in specific languages is strongly influenced by geopolitical events and national security priorities. The clearest documented case is Arabic NLP, which experienced a dramatic surge in funding following September 11, 2001. Darwish et al. [habash2022panoramic] documented the fact that “Arabic NLP gained increasing importance in the Western world especially after September 11. The USA funded large projects for companies and research centers to develop NLP tools for Arabic and its dialects”. This investment wave (2001-2010) produced fundamental resources for machine translation, speech recognition, named entity recognition, and information extraction, that continue to underpin Arabic multimodal research today.
A similar pattern is emerging for Ukrainian. Prior to February 2022, Ukrainian was a moderately-resourced Slavic language, receiving limited attention in NLP research. The Russian invasion triggered rapid mobilization: the CLARIN Knowledge Centre for Ukrainian NLP (UkrNLP-Corpora) was established in 2023 [clarin2023ukrainian], the Ukrainian Natural Language Processing Workshop (UNLP) expanded to four editions by 2025, and researchers developed numerous datasets for disinformation detection, sentiment analysis, and propaganda identification on Ukrainian social media [unlp2025proceedings]. This research is explicitly framed around information warfare: detecting “manipulative narratives” and “rhetorical manipulation techniques used to influence Ukrainian Telegram users” [akhynko2025hidden]. The geopolitical urgency has attracted Western funding and research attention that Ukrainian might not otherwise have received.
The US-China technology competition further illustrates how strategic rivalries shape NLP investment trajectories. China invested \$125 billion in AI in 2025, representing 38% of global AI investment, with NLP receiving 11% of this allocation [secondtalent2025chinese]. Chinese companies, including Baidu, Alibaba, and Tencent, are developing competitive large language models (Qwen, Yi, DeepSeek) trained on massive Chinese corpora, partly driven by the US export controls on advanced semiconductors [science2025deepseek]. This competition benefits Mandarin Chinese language resources, but does not extend to minority languages within China (Tibetan, Uyghur, Mongolian), which remain severely underrepresented despite large speaker populations. This indicates that geopolitical attention flows to languages of strategic interest to major powers, not necessarily to the most linguistically marginalized communities.
The patterns identified above reveal a troubling dynamic for truly underrepresented languages: research investment follows geopolitical salience rather than linguistic need. Languages become “high-resource” when powerful states perceive strategic value in processing them for intelligence gathering, countering disinformation, or economic competition. The 42 languages in our survey with only a single study lack this geopolitical visibility. For researchers working on such languages, this suggests that framing research around emerging strategic concerns (e.g. climate migration, regional stability, pandemic communication) may attract funding that purely linguistic motivations cannot.
Table 2: Comparison of our survey with related work on LLMs and LMMs across different focuses, modalities, languages, techniques, and additional coverage.
Survey Focus Multimodality Languages Techniques Additional Coverage
Data Creation
Fusion
Visual Enh.
Transfer
Adaptation
Gu et al. [gu2024survey] LLM-as-a-judge High-resource Evaluation, reliability, applications
Joshi et al. [joshi2020state] Language resources Low-resource Digital divide
Paullada et al. [Paullada2020DataAI] Dataset development Low-resource Data challenges
Ruder et al. [Ruder2021XTREMERTM] Cross-lingual NLP Low-resource Benchmarking
Zhao et al. [Zhao2023ASO] LLMs Both Pre-training, adaptation, utilization
Zhu et al. [Zhu2023ExtrapolatingLL] Multilingual LLM Both Cross-lingual transfer
Gan et al. [Gan2022VisionLanguagePB] Vision-language High-resource Pre-training objectives
Wang et al. [2023-Wang-survey] Pre-training High-resource Model architectures
Li et al. [Art_Large_Vision] Large VLMs High-resource Benchmark evaluations, challenges
Yin et al. [2023-yin] LMM architectures High-resource Training strategies
Xie et al. [xie2024large] Large multimodal agents High-resource Agentic AI, evaluation methods
Xu et al. [xu2024survey] Resource-efficient models Both Efficient algorithms, system designs
Alam et al. [2024-alam] LLMs for LR contexts Low-resource Capabilities, prompting, evaluation
Mu et al. [DBLP:journals/corr/abs-2503-07137] Mixture-of-Experts Both Algorithms, theory, applications
Our Survey LMMs for LR languages Low-resource Systematic taxonomy, 96 languages, 117 studies
@@
Relation to other surveys and academic contributions. Some recent surveys have explored various aspects of multimodal language models. Zhao et al. [Zhao2023ASO] provided a comprehensive overview of LMM architectures, training strategies, and applications, while Wang et al. [2023-Wang-survey] focused on pre-training techniques and model architectures. Additional surveys have examined related areas, including LLMs [gu2024survey, joshi2020state, Paullada2020DataAI, Ruder2021XTREMERTM, Gan2022VisionLanguagePB, Art_Large_Vision, xie2024large, xu2024survey, DBLP:journals/corr/abs-2503-07137], but they did not specifically address the unique challenges and solutions for LR languages in multimodal contexts. Alam et al. [2024-alam] explored LLMs for low-resource languages in multilingual, multimodal and dialectal settings, but they focused primarily on the capabilities of LLMs rather than presenting a comprehensive survey of techniques. To the best of our knowledge, our survey is the first to focus on multimodal learning for understudied languages.
As shown in Table 2, our work differs from previous surveys by specifically focusing on the intersection of multimodality and low-resource languages, while addressing all major techniques relevant to this domain. While several prior surveys have separately explored multimodality or low-resource languages, none has comprehensively examined both aspects across such a diverse range of languages and approaches.
@@
  • 1.
    We provide the first comprehensive analysis of LMMs specifically focused on LR languages, examining 117 studies across 96 languages.
  • 2.
    We develop a novel taxonomy (see Figure 3) that categorizes existing approaches into six main categories: multimodal data creation, synthetic data generation, multimodal fusion techniques, visual enhancement techniques, cross-modal transfer learning and architectural innovations.
  • 3.
@@
Organization. The remainder of this survey is organized as follows. In Section 2, we present an overview of the constructed taxonomy and discuss research trends between 2018 and 2025. Sections 3 and 4 are dedicated to resource-oriented contributions. In Section 3, we identify and categorize common approaches for dataset creation for LR languages. In Section 4, we discuss automated data generation techniques. Sections 5, 6, 7 and 8 are dedicated to method-oriented contributions. In Section 5, we categorize and compare strategies used to fuse multiple modalities. In Section 6, we discuss techniques used to enhance machine translation by using visual information. In Section 7, we present methods that perform cross-modal transfer learning. In Section 8, we analyze architectural contributions. In Section 9, we discuss current evaluation challenges and propose several ways to address the identified challenges. In Section 10, we draw our conclusions, point out current research gaps, and propose ways to mitigate them in future work.
@@ 2 Taxonomy
To organize the diverse approaches in the rapidly evolving field of LMMs for low-resource languages, we develop a comprehensive taxonomy through a systematic analysis of the 117 papers in our survey. Our methodology involves initial coding of each paper’s primary contributions and techniques, iterative refinement through thematic analysis to identify recurring patterns, and hierarchical organization of approaches based on their functional relationships and chronological development in the field. Our analysis reveals that researchers addressing the challenges of LMMs for low-resource languages typically follow a progression from resource development to architectural refinement. This progression is reflected in our taxonomy, which organizes approaches into six main categories that represent both the current state of the field and the primary research strategies for addressing challenges in the context of underrepresented languages.
In Figure 3, we systematically organize LMMs for LR languages into six main categories. The first two categories focus on constructing high-quality resources. While the first category discusses multimodal data creation either from scratch or via extending existing datasets, the second approach centers on synthetic data generation, which automatically expands available resources via back-translation and image-based generation. Building upon this work, we present several multimodal fusion techniques and provide various strategies for effectively combining this information, ranging from early and late fusion to more complex hybrid approaches. In the fourth category, we illustrate visual enhancement techniques that harness visual information through image-guided translation and visual disambiguation methods, highlighting their importance for improving translation quality and resolving ambiguities. Expanding from the single-modality solutions, the next category focuses on cross-modal transfer learning approaches that can facilitate knowledge sharing based on both modality transfer and language transfer. Finally, our last category comprises architectural innovations specifically tailored for multimodal tasks in the context of LR languages.
It is important to note that several studies naturally span multiple areas. For instance, some studies that fuse visual and textual features could also be viewed as cross-modal transfer when they leverage pre-trained vision-language models. Similarly, papers that introduce new datasets may incorporate synthetic augmentation, and architecture-focused contributions may sometimes rely on transfer-learning mechanisms. In such cases, we assign each study to the category that reflects its primary technical contribution. This principle helps maintain clear boundaries, while acknowledging the natural overlap across multimodal methods for low-resource languages.
Furthermore, to understand how these categories have evolved over time, we analyze the publication trends from 2018 to 2025. Figure 4 shows the number of papers per year in each category, illustrating how early work focused primarily on data creation and synthetic augmentation, while recent years saw an increase in fusion strategies, cross-modal transfer, and architectural innovations, reflecting the shift toward foundation-model-based approaches.
Figure 4: Number of LMM papers for LR languages published per year (2018-2025), categorized by technique: Multimodal Data Creation, Synthetic Data Generation, Multimodal Fusion Techniques, Visual Enhancement Techniques, Cross-Modal Transfer Learning, and Architectural Innovations. Best viewed in color.
Our taxonomic structure not only organizes existing research, but also highlights the inter-dependencies between different approaches and reveals gaps in current research, particularly in the exploration of complex multimodal combinations involving video and speech for low-resource languages. We structure the remainder of this article according to our novel taxonomy shown in Figure 3.
@@ 3 Multimodal Data Creation
Dataset creation from scratch. Dataset creation from scratch has emerged as a crucial approach for enabling multimodal research in LR languages, particularly for sentiment analysis and specific language tasks. Multiple research teams have focused on creating specialized datasets through direct data collection and annotation, such as collecting Arabic videos with multimodal features for sentiment analysis [2018-najadat], building comprehensive Tamil and Malayalam video review datasets [2021-Chakravarthi], developing new corpora for languages such as Malay [2024-taylor], creating speech translation resources for Fongbe [kponou-etal-2024-ffstc] and compiling Arabic multimodal sentiment collections [2023-haouhat]. A significant trend has been the creation of meme-based datasets, with efforts focused on Bengali, through MemoSen and MUTE [2022-hossain, 2022-hossain-mute], and Romanian, through RoMemes [Pais2024RoMemesAM], all incorporating multiple levels of annotation.
@@
These dataset creation efforts have expanded beyond sentiment analysis to encompass other crucial applications, such as sign language recognition with ArabSign [2022-hamzah], and multi-purpose datasets like BIG-C for Bemba [2023-sikasote]. Additionally, the creation of a Manipuri-English parallel corpus with accompanying audio recordings for speech-to-text translation [2021-sanayai] provides an important resource for research in low-resource languages. More recently, Farsi et al. [farsi-etal-2025-persian] introduced a comprehensive suite of multimodal datasets for Persian, covering tasks such as VQA, OCR, visual abstraction reasoning, and cultural knowledge grounding. These projects typically involve careful quality control by using multiple annotators, standardized recording environments, and expert validation, demonstrating a shift toward building comprehensive resources specifically designed for LR languages, rather than relying on translation or transfer from high-resource languages.
@@
Dataset extension. In addition to building data from scratch in the context of LR language understanding, there have been several efforts for leveraging existing datasets of rich-resource languages and building upon them. Sen et al. [2022-sen] introduced the Bengali Visual Genome (BVG) dataset, which extends the Visual Genome dataset [krishna2017visual] with Bengali translations and annotations, enabling the development and evaluation of multimodal models for Bengali-English machine translation (MT) and image captioning. Similarly, Abdulmumin et al. [2022-abdulmumin] created the Hausa Visual Genome (HaVG) dataset by translating a subset of the Visual Genome dataset into Hausa, providing a valuable resource for English-to-Hausa multimodal MT. Building upon prior work and continuing the focus on the Hausa language, Parida et al. [2023-parida] introduced the Hausa Visual Question Answering (HaVQA) dataset, which adapts question-answer pairs from the Visual Genome dataset to the Hausa language through manual translation, creating the first visual question-answering (VQA) dataset for Hausa. Extending this trend to Indian languages, Parida et al. [parida-etal-2025-ovqa] introduced OVQA, a multimodal dataset for the Odia language, by translating over 6,000 question-answer pairs and associated captions from the Visual Genome dataset into Odia. Similarly, Anwar et al. [2023-anwar] introduced MuAViC, a multilingual audio-visual corpus providing 1,200 hours of audio-visual speech across 9 languages, establishing the first open benchmark for audio-visual speech-to-text translation.
@@
Apart from the focus on African languages, Saichyshyna et al. [2023-saichyshyna] extended the Multi30K dataset [elliott2016multi30k] to include Ukrainian translations and captions, facilitating integrated vision and language research in Ukrainian. More recently, Lovenia et al. [2024-lovenia] presented SEACrowd, a comprehensive multilingual and multimodal data hub and benchmark suite for Southeast Asian languages, which covers 13 tasks across three modalities (text, image, and audio) and 38 Southeast Asian indigenous languages, while Lent et al. [lent-etal-2024-creoleval] introduced CreoleVal, an extensive collection of benchmarks for 28 Creole languages, addressing the significant resource gap for these historically marginalized language varieties.